# RILe: Reinforced Imitation Learning

### v2

@@ 5.1 Evolving Reward Function
To evaluate the impact of RILe’s trainer agent on the learning process compared to existing methods, we designed a maze experiment. The task is to navigate from a fixed start on the left side to a goal in the top-right corner, avoiding obstacles.
Using a single expert demonstration, we trained RILe, GAIL, and AIRL. Fig. 3 shows how the student’s reward function evolves during training with all three methods. For RILe, we plot the reward function learned by the trainer. For GAIL, we visualize the discriminator output, and for AIRL, the reward term under the discriminator. The columns represent reward plots at 25%, 50%, 75%, and 100% of training completion for each algorithm, respectively. We present the student agent’s trajectory from the previous epoch on reward plots to demonstrate how reward functions adapt to the student agent’s progress.
@@
Figure 3: Evolution of reward functions during training for (a) RILe, (b) GAIL, and (c) AIRL in a continuous maze environment. Columns show reward landscapes at 25%, 50%, 75%, and 100% of training completion (left to right). The expert’s trajectory is shown in red, while the student agent’s trajectory from the previous training epoch is in black. Color gradients represent reward values, with darker colors indicating lower rewards and brighter colors indicating higher rewards. Purple squares represent obstacles. RILe demonstrates a more adaptive reward function that evolves with the student’s progress, while GAIL and AIRL maintain relatively static reward landscapes throughout training.
The discriminator-based reward functions of GAIL and AIRL remain relatively static, largely independent of the student agent’s current performance. This static behavior is because the discriminators in GAIL and AIRL focus on distinguishing expert data and progressively improve throughout training. In contrast, the trainer in RILe adapts itself based on the current learning progress of the student’s policy.
As the first column shows, when the student’s trajectory does not resemble the expert, RILe’s trainer encourage exploration towards the expert path. This indicates that the trainer can provide informative rewards despite potentially low discriminator feedback. Then, in the second column, as the student progresses, RILe’s reward function adapts dynamically. When the student learns to reach the bottom-right, the trainer shifts high rewards to the top-left, guiding the agent to explore that area. Third column shows as the student approaches the goal, the trainer increases rewards around it while maintaining rewards in specific areas (e.g., the left part) to prevent the agent from getting stuck. Finally, as the fourth column shows, the trainer focuses high rewards on the top-right area near the goal while maintaining a path of rewards that guides the student along the expert trajectory.
RILe’s evolving reward function demonstrates its ability to provide meaningful guidance even when the discriminator easily identifies non-expert policies. By adapting to the student’s current capabilities, RILe creates a dynamic learning curriculum that encourages exploration and gradual improvement towards expert-like behavior. Our results demonstrate that the trainer effectively associates rewards with future behavior improvements, even if delayed. It provides incremental rewards for actions moving the student in the right direction, and adapts the reward function to match the student’s abilities, gradually guiding them toward expert behavior.
5.2 Motion-Capture Data Imitation for Robotic Continuous Control
We evaluate RILe’s performance on the LocoMujoco benchmark, which presents a challenging task of imitating motion-capture (MoCap) data for various robotic control tasks. This benchmark is particularly demanding due to its high dimensionality and the absence of action data in the MoCap recordings which precludes the use of methods like Behavioral Cloning and IQLearn that require complete state-action pairs.
Table 1: Test and validation results on seven LocoMujoco tasks.
Test Seeds Validation Seeds
RILe GAIL AIRL RILe GAIL AIRL Expert
Atlas Walk 870.6 792.7 300.5 895.4 918.6 356 1000
Atlas Carry 850.8 669.3 256.4 889.7 974.2 271.9 1000
Talos Walk 842.5 442.3 102.1 884.7 675.5 103.4 1000
Talos Carry 220.1 186.3 134.2 503.3 338.5 74.1 1000
UnitreeH1 Walk 966.2 950.2 568.1 980.7 965.1 716.2 1000
UnitreeH1 Carry 788.3 634.6 130.5 850.6 637.4 140.9 1000
Humanoid Walk 831.3 181.4 80.1 970.3 216.2 78.2 1000
Table 1 presents the results for seven LocoMujoco tasks, comparing RILe with GAIL and AIRL across both test and validation seeds. The expert performance, set at 1000 for all tasks, serves as the upper benchmark. RILe demonstrates superior performance in all scenarios, particularly excelling in generalization to new initial conditions as evidenced by the test seed results.
5.3 Impact of Expert Data on Trainer-Student Dynamics
@@
Figure 4: Impact of explicit usage of expert data on RILe’s performance. Red markers and yellow markers show normalized scores and normalized steps, respectively. Square markers show IQ-Learn and BC results for comparisons. The percentage of expert data in the buffers are presented around the figure. More expert data in buffers speeds convergence but can reduce final performance, especially at 100%, showing how the explicit usage of excess expert data hinders trainer-student adaptation in RILe
We study how explicitly incorporating expert data into RILe’s training affects the trainer’s ability to adapt to the student’s needs, analyzing performance and convergence time when expert data is added to the agents’ replay buffers. For this experiment, we used MuJoCo’s Humanoid environment [22, 21] and a single expert trajectory from [24]. We varied the proportion of expert data in the replay buffers from 0% to 100%; for example, 25% means a quarter of the buffer is expert data and 75% is from agent interactions.
As shown in Fig. 4, introducing expert data led to faster convergence but decreased performance. Notably, when environmental interactions were completely replaced by the expert data (100% case), the student’s performance declined significantly. Excessive expert data appears to hinder the trainer’s ability to adapt to the student, disrupting RILe’s dynamic learning process. For comparison, we include results from IQ-Learn and Behavioral Cloning (BC), which rely explicitly on expert data. Despite their exclusive use of expert data, neither method matched RILe’s performance, even when RILe used substantial amounts of expert data.
Our findings underscore the importance of balancing environmental interactions and expert demonstrations in buffers in RILe. While some expert data accelerates learning, too much impedes the trainer’s adaptability, underscoring the unique strength of RILe’s interactive learning approach.
@@ 6 Discussion
Our experiments have demonstrated that RILe consistently outperforms baseline models across various settings, surpassing state-of-the-art adversarial IL/IRL methods. RILe’s key strength lies in its adaptive learning approach, where the trainer agent dynamically adjusts the reward function based on the student’s current capabilities. This enables more efficient learning, particularly in complex environments.
However, RILe faces challenges, primarily in maintaining policy stability with a changing reward function. Our solution of freezing the trainer after the student achieves half the expert score is effective but potentially limiting. Additionally, the discriminator’s tendency to overfit quickly presents another significant challenge. Future work should focus on establishing bounds for reward agent updates to enhance stability without compromising adaptability, developing a stronger theoretical understanding of RILe’s convergence properties, and exploring discriminator-less approaches.
Despite these challenges, RILe has demonstrated advantages in adaptability, robustness, and generalization, thereby making it a promising direction for future research in imitation learning and opening up new possibilities for dynamic and responsive learning processes.
Appendix A Justification of RILe
@@
A.1 Lemma 2:
@@
Proof for Lemma 2: The training objective for the discriminator is framed as a binary classification problem over mini-batches from expert demonstrations and student-generated trajectories. The discriminator’s loss function $\mathcal{L}_{D}(\phi)$ is the binary cross-entropy loss, which for a mini-batch of size $n$ is defined as:
$\mathcal{L}_{D}(\phi)=-\frac{1}{n}\sum_{i=1}^{n}\Big{[}y_{i}\log(D_{\phi}(s_{i},a_{i}))+(1-y_{i})\log(1-D_{\phi}(s_{i},a_{i}))\Big{]},$ (16)
where $(s_{i},a_{i})$ are sampled state-action pairs from the combined replay buffer $\mathcal{D}=\mathcal{D}_{S}\cup\mathcal{D}_{E}$, with corresponding labels $y_{i}$ indicating whether the pair is from the expert $(y_{i}=1)$ or the student $(y_{i}=0)$. The stochastic gradient descent update rule for optimizing $\mathcal{L}_{D}(\phi)$ is then given by:
$\phi\leftarrow\phi-\eta_{D}\nabla_{\phi}\mathcal{L}_{D}(\phi),$ (17)
where $\eta_{D}$ is the learning rate for the discriminator.
Through iterative updates, $D_{\phi}(s,a)$ will converge to $P(\pi_{E}|s,a)$, provided the minimization of $\mathcal{L}_{D}(\phi)$ progresses according to the theoretical foundations of stochastic gradient descent. The convergence relies on the assumption that the student’s policy $\pi^{\theta_{S}}$ and the expert policy $\pi_{E}$ induce stationary distributions of state-action pairs, such that the discriminator’s data source is consistently representing both policies over time.
By minimizing $\mathcal{L}_{D}(\phi)$, we seek $\phi^{*}$ such that:
$\phi^{*}=\arg\min_{\phi}\mathcal{L}_{D}(\phi).$ (18)
Under typical conditions for convergence in stochastic gradient descent, the convergence to a local minimum or saddle point can be guaranteed. The discriminator’s ability to distinguish between student and expert pairs improves as $\mathcal{L}_{D}(\phi)$ is minimized, implying that $\lim_{n_{batch}\rightarrow\infty}D{\phi^{*}}(s,a)=P(\pi_{E}|s,a)$, where $n_{batch}$ is the number of batches.
Appendix B Additional Experiments
B.1 Noisy Expert Data
(a) Expert traj.
(b) RILe traj.
(c) GAIL traj.
(d) AIRL traj.
(e) IQLearn traj.
(f) RILe val.
(g) GAIL val.
(h) AIRL val.
(i) IQLearn val.
Figure 5: In a 5x5 grid environment with lava, (a) the expert trajectory is characterized by noisy data that passes through lava without resulting in death. (c) GAIL, (d) AIRL and (e) IQLearn learn to imitate the expert’s path precisely, leading them to either get stuck near the lava or enter it and perish. (b) RILe avoids the noisy data, better mimics the expert in later stages, and successfully reaches the goal. Subfigures (f-i) display the value tables for RILe, GAIL, AIRL, and IQLearn respectively. The optimal path, derived from the reward of the trainer or discriminator, is highlighted with green lines.
To demonstrate the advantage of using RL to learn the reward function in RILe, as opposed to deriving the reward directly from the discriminator in AIL and AIRL, we designed a 5x5 MiniGrid experiment. The grid consists of 4 lava tiles that immediately kill the agent if it steps in it, representing terminal conditions. The goal condition of the environment is reaching the green tile.
The expert demonstrations are imperfect, depicting an expert that passes through a lava tile without being killed and still reaches the green goal tile. Using this data, we trained the adversarial approaches with a perfect discriminator, which provides a reward of $0.99$ if the visited state-action pair stems from the expert and $0.01$ otherwise. These values were chosen over 1 and 0 because both AIRL and GAIL use the logarithm of the discriminator output to calculate rewards.
Results are presented in Fig. 5. The value graphs (Fig. 5e-g) are attained by computing the value of each grid cell $c_{i}$ as $\sum_{a\in A}D(c_{i},a)$ for AIRL and GAIL, and $\sum_{a\in A}\pi_{T}(c_{i},a)$ for RILe. Fig. 5a shows the expert trajectory.
GAIL (Fig. 5c), AIRL (Fig. 5d) and IQLearn (Fig. 5e) fail to reach the goal, as their agents either become stuck or are directed into lava.
In contrast, RILe (Fig. 5d) successfully reaches the goal, demonstrating its ability to navigate around imperfections in expert data. The difference in the value graphs between RILe and the baselines intuitively explains this outcome. In AIL and AIRL (Fig. 5f-g), the optimal paths, defined by the actions most rewarded by their discriminators, follow the noisy expert data perfectly. Similarly, in IQLearn, the agent tries to match expert state-actions as closely as possible, minimizing any deviation from the expert trajectory. In contrast, RILe’s trainer agent, trained using RL, adds an extra degree of freedom in the adversarial IL/IRL setting. By providing rewards that maximize cumulative returns from the discriminator, rather than deriving the reward directly from its output, the value graph (Fig. 5f) can learn to circumvent the lava tile in order to follow the expert trajectory to the goal. Consequently, the optimal path of the student agent can overcome the sub-optimal state suggested by the noisy expert demonstration. Since the student agent is guided by the trainer to also match the expert trajectory, it remains close to this path after passing the lava tiles.
B.2 Robustness to Noise in the Expert Data

### v3

@@ 5.1 Evolving Reward Function
To answer the first question about RILe’s reward-learning strategy, we compare RILe’s performance with AI(R)L baselines in a maze setting. In this environment, the agent must navigate from a fixed start to a goal while avoiding static obstacles; we use a single expert demonstration.
@@
Figure 3: Reward Function Comparison. Evolution of reward functions during training for (a) RILe, (b) GAIL, and (c) AIRL in a continuous maze environment. Columns show reward landscapes at 25%, 50%, 75%, and 100% of training completion (left to right). The expert’s trajectory is shown in red, while the student agent’s trajectory from the previous training epoch is in black. Color gradients represent reward values, with darker colors indicating lower rewards and brighter colors indicating higher rewards. Purple squares represent obstacles. RILe demonstrates a dynamic reward function that adapts with the student’s progress, while GAIL and AIRL maintain relatively static reward landscapes throughout training.
Figure 3 shows how each method’s learned reward function evolves during training. For RILe, we plot the trainer’s learned reward function. For GAIL and AIRL, we visualize the discriminator outputs. The columns represent reward landscapes at 25%, 50%, 75%, and 100% of the total training process, and each subplot overlays the student’s trajectory from the previous epoch.
RILe’s reward function dynamically adapts to the student’s current policy, providing guidance that encourage suboptimal actions which eventually lead the student closer to the expert trajectory. By contrast, GAIL and AIRL’s reward functions remain relatively static. Specifically, the first column in Figure 3 shows RILe’s trainer encouraging exploration toward the bottom-right of the maze, which is initially suboptimal but helpful in the long run. As the student learns to reach the lower part of the maze, RILe shifts high-reward regions toward the top-left (second column), again encouraging incremental progress. The third column illustrates how RILe boosts rewards near the goal while still maintaining some incentive around other key areas to keep the agent from getting stuck. Overall, RILe’s evolving reward function serves as a curriculum that promotes gradual improvement toward expert-like performance. This dynamic reward adaptation gets important in higher-dimensional tasks as we show in Section 5.3.
5.2 Reward Function Dynamics
To address the second question on adaptability, we quantify how the reward function evolves and its relationship to the student’s performance. We compare RILe with GAIL, DRAIL-GAIL, and DRAIL-RILe in a high-dimensional robotic control scenario (learning to walk with UnitreeH1 robot).
We introduce three metrics (see Appendix D.2 for details): (1) RFDC (Reward Function Distribution Change): Wasserstein distance between reward distributions over consecutive training intervals, capturing overall shifts in reward space, (2) FS-RFDC (Fixed-State Reward Function Distribution Change): Mean absolute deviation of reward values at a fixed set of expert states over time, (3) CPR (Correlation between Performance and Reward): Pearson correlation between changes in the reward function and changes in the student’s performance.
(a) RFDC
(b) FS-RFDC
(c) CPR
Figure 4: Dynamics of Reward Functions. (a) Reward Function Distribution Change (RFDC): Wasserstein distance between reward function distributions. (b) Fixed-State Reward Function Distribution Change (FS-RFDC): Mean absolute deviation of reward values for a fixed set of expert states. (c) Correlation between Performance and Reward (CPR): Pearson correlation between changes in the reward function and changes in the student’s performance.
5.2.1 Adaptability of the Learned Reward Function
We assess how adaptive the reward function learned by RILe is compared to that of AIL. Fig. 4(a) presents changes in reward distributions over 10,000 consecutive steps. RILe exhibits the highest adaptability in its reward function, aligning with our goal of having the reward function adapt based on the student’s learning stage. The advanced discriminator in DRAIL reduces the need for drastic reward function changes, yet RILe remains more adaptive than GAIL. Since the changing student policy indirectly affects the Fig. 4(a), we also show changes in reward values for the fixed set of states in Fig. 4(b). Again, RILe’s reward function is the most adaptive among all methods.
Higher adaptability of RILe ensures that the reward signal remains aligned with incremental performance improvements, enabling the student to receive more timely and effective guidance throughout training.
5.2.2 Correlation between the Learned Reward and the Student Performance
We evaluate how changes in the reward function correlate with improvements in student performance. To this end, Fig. 4(c) presents the Pearson correlation between student’s performance and reward updates. DRAIL-RILe achieves the highest positive correlation, indicating that it learns the most effective rewards for improving student performance. RILe ranks second, demonstrating that the trainer agent effectively helps the student achieve better scores even with the help of a naive-discriminator. In contrast, GAIL’s correlation starts positive but soon turns negative and remains so throughout training. We hypothesize that this occurs because the discriminator in GAIL tends to saturate as training progresses. While the discriminator’s reward signal effectively guides learning early on, its increasingly static nature at later stages fails to capture subtle performance improvements, leading to a negative correlation.
5.3 Imitation in Continuous Control Tasks
5.3.1 Motion-Capture Data Imitation for Robotic Continuous Control
To answer the third question, we first test RILe in high-dimensional robotic locomotion tasks [22], where the agent must imitate motion-capture data for various robotic bodies. This benchmark is especially demanding due to its complexity and dimensionality.
Table 1: Test results on seven LocoMujoco tasks.
RILe GAIL AIRL IQ BCO GAIfO DRAIL GAIL DRAIL RILe Expert
Walk Atlas 870.6 792.7 300.5 30.9 21.0 834.2 834.4 899.1 1000
Talos 842.5 442.3 102.1 4.5 11.9 710.0 787.7 896.6 1000
UnitreeH1 966.2 950.2 568.1 8.8 34.8 526.8 940.8 995.8 1000
Humanoid 831.3 181.4 80.1 4.5 3.5 706.5 814.6 527.6 1000
Carry Atlas 850.8 669.3 256.4 36.8 20.3 810.1 516.6 317.1 1000
Talos 220.1 186.3 134.2 10.5 10.3 212.5 836.7 840.5 1000
UnitreeH1 788.3 634.6 130.5 14.4 21.1 604.5 796.7 909.5 1000
Table 1 presents results for seven LocoMujoco tasks across different test seeds (see Appendix D.3 for details). Overall, these results underscore that RILe outperforms the AIL/IL/IRL baselines, and benefits even more from the enhancements provided by the DRAIL variants, achieving performance levels close to the expert. The performance variations across tasks indicate that although RILe’s adaptive reward function is highly effective, task-specific factors also play a role, suggesting avenues for future research. This overall performance aligns with our claim that an adaptive reward function is crucial for mastering complex, high-dimensional behaviors
Table 2: Test results on four MuJoCo tasks.
RILe GAIL AIRL IQ
Humanoid 5928 5709 5623 327
Walker2d 4435 4906 4823 270
Hopper 3417 3361 3014 310
HalfCheetah 5205 4173 3991 755
@@ 5.3.1 Motion-Capture Data Imitation for Robotic Continuous Control
Figure 5: Explicit Usage of Expert Data. Red and yellow markers show normalized scores and steps, respectively. Expert data usage speeds the training of RILe but reduces final performance.
5.3.2 Learning from Demonstrations
We further evaluate RILe’s performance on four standard MuJoCo tasks [23, 24], Humanoid, Walker2d, Hopper, and HalfCheetah, with baseline comparisons widely reported in prior work.
Table 2 summarizes test results. Overall, these results demonstrate that RILe consistently achieves competitive or superior performance compared to established AIL/AIRL/IRL methods. In particular, on high-dimensional tasks such as Humanoid, RILe’s performance advantage is evident, underscoring the effectiveness of its adaptive reward function in handling complex continuous control challenges.
5.4 Explicit Usage of Expert Data
To answer the fourth question, we use the Humanoid environment in MuJoCo [23, 24] with a single expert trajectory from [7]. We vary the proportion of expert data in the replay buffers from 0% to 100%. For instance, 25% indicates a quarter of each buffer is expert data and 75% is agent data (see Appendix D.5 for details).
Figure 5 illustrates how changing the proportion of expert data in both the trainer’s and the student’s replay buffers influences RILe’s training. Introducing more expert data accelerates RILe’s convergence (needs fewer training steps), but reduce the final performance. In the extreme case of 100% expert data, the student’s performance drops substantially. This indicates that too much expert data hamper the trainer’s ability to adapt to the student’s real-time needs, disrupting RILe’s context-sensitive reward customization.
We also compare RILe’s results with IQ-Learn and BC, both of which rely heavily on expert data. Even when RILe is trained with substantial amounts of expert data, it still generally performs better, indicating that RILe’s adaptive reward-shaping provides a crucial edge over supervised methods.
@@ 6 Discussion
As our experiments demonstrate, RILe consistently outperforms baseline models across various tasks, thanks to its adaptive learning approach, where the trainer agent continuously adjusts the reward based on the student’s current learning stage.
Our maze experiments provide an interpretable example of how the trainer agent tailors its rewards. By encouraging actions that might seem suboptimal for immediate imitation but advantageous for long-term learning, RILe establishes a curriculum that ultimately boosts performance. This adaptive strategy helps RILe to achieve superior results in our continuous control experiments, where reward shaping becomes especially critical in high-dimensional settings.
Nonetheless, policy stability remains challenging when the reward function is constantly evolving. Freezing the trainer (see Appendix C) stabilizes learning but halts further adaptation, and the discriminator itself tends to overfit quickly. Future work could explore fully cooperative multi-agent RL techniques to allow ongoing adaptation, investigate ways to bound or regulate trainer updates, and consider discriminator-less formulations for reward learning.
Despite these challenges, RILe shows that cooperatively learning the policy and the reward function can offer significant advantages over static or iteratively updated methods. By providing dynamic and tailored rewards, RILe effectively guides the student through complex tasks. We believe this opens up new possibilities for responsive and adaptive learning frameworks in imitation learning and beyond.
Appendix A POMDP of the Trainer
Partially Observable Markov Decision Process (POMDP) of the trainer is defined as $\text{POMDP}_{T}=(S_{T},A_{T},\Omega_{T},T_{T},O_{T},R_{T},\gamma)$. Here, $T_{T}=\{P(.\mid f^{T},a^{T})\}$ is the transition dynamics where $P(.\mid f^{T},a^{T})$ is the state distribution upon taking action $a\in A_{T}$ in state $f\in S_{T}$. The transition function incorporates the student’s policy $\pi_{S}$, which evolves in response to the rewards provided, reflecting the hidden dynamics due to the unobserved $\pi_{S}$. The observation function $O_{T}=\{P(s^{T}\mid f^{T},a^{T})\}$ defines the probability of observing $s^{T}\in\Omega_{T}$ given the state $(f^{T},a^{T})$. The trainer deterministically observes the student’s state-action pair, so $P(s^{T}=(s^{S},a^{S})\mid f^{T},a^{T})=1$, where $f^{T}=(s^{S},a^{S},\pi_{S})$.
Appendix B Justification of RILe
@@
B.1 Lemma 1:
Given the discriminator $D_{\phi}$, the trainer agent optimizes its policy $\pi^{\theta_{T}}$ via policy gradients to provide rewards that guide the student agent to match expert’s state-action distributions.
Proof for Lemma 1 The student agent, $\pi_{S}(a^{S}_{t}|s^{S}_{t})$, interacts with the environment and generates state-action pairs as $(s^{S}_{t},a^{S}_{t})$. The trainer agent observes these pairs and provides a reward $r^{S}_{t}=a^{T}_{t}=\pi_{T}(a^{T}_{t}|(s^{S}_{t},a^{S}_{t}))$ to the student, where $a^{T}_{t}\in[-1,1]$ is the trainer’s action. We have $D_{\phi}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ as the discriminator, parameterized by $\phi$, which outputs the likelihood that a given state-action pair $(s,a)$ originates from the expert, as opposed to the student.
The trainer’s reward at timestep $t$ is:
$r_{t}^{T}=\upsilon(D_{\phi}(s_{t}^{T}))a_{t}^{T}$ (11)
where $s_{t}^{T}=(s^{S}_{t},a^{S}_{t})$ is the trainer’s observation, $D_{\phi}(s_{t}^{T})$ is the discrimantor output that estimates the likelihood that $s_{t}^{T}$ comes from the expert data, and $\upsilon(D)=2D-1$ is a scaling function that maps discriminator’s output to the range $[-1,1]$.
The trainer maximizes the expected cumulative reward:
$J_{T}(\pi_{T})=\mathbb{E}_{\pi_{T},\pi_{S}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}^{T}\right]$ (12)
where $\gamma\in[0,1)$ is the discount factor. In other words, trainer aims to find the policy that maximizes $J_{T}(\pi_{T})$: $\pi^{*T}=\operatorname*{arg\,max}_{\pi^{T}}J_{T}(\pi_{T})$.
From the policy gradient theorem, the gradient of the trainer’s objective with respect to the policy parameters, $\theta_{T}$, is:
$\nabla_{\theta_{T}}J_{T}(\pi_{T})=\mathbb{E}_{\pi_{T},\pi_{S}}\left[\nabla_{\theta_{T}}\log\pi_{T}(a_{t}^{T}|s_{t}^{T})Q_{T}(s_{t}^{T},a_{t}^{T})\right]\$ (13)
where $Q_{T}(s_{t}^{T},a_{t}^{T})$ is the action-value function of the trainer. The action-value function, $Q_{T}(s_{t}^{T},a_{t}^{T})$, and the value function, $V_{T}(s^{T}_{t})$ is defined by Bellman equation as:
$Q_{T}(s_{t}^{T},a_{t}^{T})$ $=r_{t}^{T}+\gamma\mathbb{E}_{s^{T}_{t+1}}\left[V_{T}(s^{T}_{t+1})\right]$ (14)
$V_{T}(s^{T}_{t+1})$ $=\mathbb{E}_{a_{t}^{T}\sim\pi_{T}}\left[Q_{T}(s_{t}^{T},a_{t}^{T}))\right]$ (15)
The trainer aims to maximize $Q_{T}(s^{T}_{t},a^{T}_{t})$ to satisfy Equation 13. Since $r^{T}_{t}$ depends directly on $D_{\phi}(s^{T}_{t})$ and $a^{T}_{t}$, the trainer learns to select $a^{T}_{t}$ that maximizes $Q_{T}(s^{T}_{t},a^{T}_{t})$ Considering that $a^{T}_{t}\in[-1,1]$, the immediate reward $r^{T}_{t}$ is maximized when $a^{T}_{t}$ has the same sign as $\upsilon(D_{\phi}(s^{T}_{t}))$ Therefore, the optimal action $a^{*T}_{t}$ is:
$\alpha^{*T}_{t}=\begin{cases}1,&\text{if }D\phi(s_{t}^{T})>0.5,\\ -1,&\text{if }D\phi(s_{t}^{T})<0.5,\\ \text{any value in }[-1,1],&\text{if }D\phi(s_{t}^{T})=0.5.\end{cases}$ (16)
Equation 16 implies the trainer assigns positive rewards to student state-action pairs that the discriminator assesses as more likely to be from the expert ($D_{\phi}(s^{T}_{t})>0.5$) and negative rewards to those unlikely to be from the expert ($D_{\phi}(s^{T}_{t})<0.5$). By this mechanism, the trainer’s policy optimization relies on the discriminator’s assessment to assign rewards that encourage expert-like behavior.
All in all, the derivative of the trainer’s expected reward, Equation 13, with respect to its policy parameters is rewritten as:
$\nabla_{\theta_{T}}J_{T}(\pi_{T})=\mathbb{E}_{\pi_{T},\pi_{S}}\left[\nabla_{\theta_{T}}\log\pi_{T}(a_{t}^{T}|s_{t}^{T})\left((2D_{\phi}(s_{t}^{T})-1)a_{t}^{T}+\gamma Q_{T}(s_{t+1}^{T},a_{t+1}^{T})\right)\right]$ (17)
The trainer adjusts $\pi_{T}$ to output high rewards when $D_{\phi}(s_{t}^{T})$ is high. Therefore the trainer learns to assign higher rewards to student behaviors that are more similar to the expert behaviors, according to the discriminator.
B.2 Lemma 2:
@@
Proof for Lemma 2: The discriminator’s objective is to distinguish between state-action pairs generated by the expert and those generated by the student. The training objective for the discriminator is framed as a binary classification problem over expert demonstrations and student-generated trajectories. The discriminator’s loss function $\mathcal{L}_{D}(\phi)$ is the binary cross-entropy loss, which is defined as:
$L_{D}(\phi)=-\mathbb{E}_{(s,a)\sim p_{E}}[\text{log}(D_{\phi}(s,a))]-\mathbb{E}_{(s,a)\sim p_{\pi_{S}}}[\text{log}(1-D_{\phi}(s,a))].$ (18)
where $p_{E}(s,a)$ is the state-action distribution of the expert policy, and $p_{\pi_{S}}(s,a)$ is the state-action distribution of the student agent. Considering that $x=(s,a)$, this loss can be rewritten as:
[…]
As presented in [25], the optimal discriminator that minimizes this loss, $D^{*}_{\phi}$, is:
$D^{*}_{\phi}(x)$ $=\frac{p_{E}(x)}{p_{E}(x)+p_{\pi_{S}}(x)},$ (21)
$D^{*}_{\phi}(s,a)$ $=\frac{p_{E}(s,a)}{p_{E}(s,a)+p_{\pi_{S}}(s,a)}.$ (22)
This shows that the optimal discriminator estimates the probability that a state-action pair comes from the expert policy, normalized by the total probability from both expert and student policies.
Appendix C Training Strategies
The introduction of the trainer agent into the AIL framework introduces instabilities that can hinder the learning process. To address these challenges, we employ three strategies.
Freezing the Trainer Agent Midway: Continuing to train the trainer agent throughout the entire process can lead to overfitting on minor fluctuations in the student’s behavior. This overfitting causes the trainer to assign inappropriate negative rewards, which diverts the student away from expert behavior—especially since the student agent may fail to interpret these subtle nuances correctly in the later stages of training. To prevent this, we freeze the trainer agent once its critic network within the actor-critic framework converges during the training process.
We consider the trainer’s critic network to have converged when the change in the exponential moving average (with a smoothing factor of 0.99) of the critic output and its variance over a window of 50000 training iterations fall below a certain threshold. In all our experiments, this threshold is set to 0.1, which we found empirically after our hyperparameter search (see Appendix H). This threshold works for all settings where the reward is bounded between $-1$ and $1$, which is the case for all our experiments.
Reducing the Trainer’s Target Network Update Frequency: We decrease the target network update frequency of the trainer agent to half that of the student agent. After our hyperparameter sweeps (see Appendix H), we empirically found that updating at half the student agent’s frequency works best. This adjustment aims to prevent overestimation bias in the trainer’s value function and to slow down its learning pace. By updating less frequently, the trainer provides more consistent and reliable reward signals. This steadier guidance helps the student agent better understand and adapt to the trainer’s rewards, facilitating more stable learning.
Increasing the Student Agent’s Exploration: We increase the exploration rate of the student agent compared to standard AIL methods. We implement an epsilon-greedy strategy within the actor-critic framework, allowing the student to occasionally take random actions. This increased exploration enables the student to visit a wider range of state-action pairs. Consequently, the trainer agent receives diverse input, helping it learn a more effective reward function. This diversity is crucial for the trainer to observe the outcomes of various actions and to guide the student more effectively toward expert behavior.
Appendix D Experimental Settings
D.1 Evolving Reward Function
We use single expert demonstration in this experiment. For RILe, we plot the reward function learned by the trainer. For GAIL, we visualize the discriminator output, and for AIRL, the reward term under the discriminator.
D.2 Reward Function Dynamics
In this experiment, we select the student agent’s hyperparameters to be identical to those used in GAIL, ensuring that the only difference between the agents is the reward function. Therefore, we use the best hyperparameters identified for GAIL, applied to both GAIL and RILe, from our hyperparameter sweeps in Appendix H.
RFDC: We calculate the Wasserstein distance between reward distributions over consecutive 10,000-step training intervals, denoted as times $t$ and $t+10,000$. This metric quantifies how much the overall reward distribution shifts over time. Changes in reward distributions depend both on the reward function and the student policy updates. Since we use the same student agent with the same hyperparameters, higher RFDC values still indicate that the reward function is adapting more dynamically in response to the student’s learning progress.
FS-RFDC: We compute the mean absolute deviation of rewards between consecutive 10,000-step training intervals for a fixed set of states derived from expert data. As the fixed set, we use all the states in the expert data. Since the states used for calculating rewards are fixed, changes in this value purely depend on the reward function updates. This metric assesses how the reward values for specific states change over time.
CPR: We evaluate how changes in the reward function correlate with improvements in student performance. We store rewards from both the learned reward function and the environment-defined rewards in separate buffers. In other words, we collect samples from two reward functions: the learned reward function and the environment-defined reward function. The environment rewards consider the agent’s velocity and stability. Every 10,000 steps, we calculate the Pearson correlation between these rewards and empty the buffers. This metric evaluates whether increases in the learned rewards relate to performance enhancements.
D.3 Motion-Capture Data Imitation for Robotic Continuous Control
During training, we use 8 different random seeds and 8 distinct initial positions for the robot. The validation setting mirrors the training conditions: we sample initial positions from the same set of 8 possibilities and use the same random seeds. In this setting, the student agent selects actions deterministically, allowing us to assess its performance under familiar conditions.
For the test setting, we evaluate the policy’s ability to generalize to new, unseen scenarios. We modify the initial positions of the robot by randomly initializing it in stable configurations not included in the fixed set used during training. Additionally, we use different random seeds from those in training, introducing new random variations that affect the environment’s dynamics during state transitions. This setup enables us to assess how well the learned policy performs when faced with novel initial conditions and environmental changes.
D.4 Learning from Demonstrations
Each method is trained using 25 expert trajectories provided in the IQ-Learn paper [7]. We use single seed for the training, and after the training, run experiments with 10 different random seeds and report the mean and standard deviation of the results.
D.5 Impact of Expert Data on Trainer-Student Dynamics
In this experiment, both seeds and initial positions in the test setting are different from the training one, and we report values from the test setting.
For every percentage of the expert-data in buffers, we continue trainings of both the trainer agent and the student agent of RILe. For instance, in 100% expert data in the trainer’s buffer case, both the student and the discriminator are trained normally using samples from the student agent. However, we didn’t include student’s state-action pairs to the trainer’s buffer, instead, we filled that buffer with a batch of expert data, and updated the trainer regularly using this modified buffer. Similarly, in 100% expert data in the student’s buffer case , we trained the trainer agent and the discriminator normally, using samples from the student agent. However, student’s state-actions pairs are not included in the student’s buffer, and student agent is updated just by using expert state-action pairs, using rewards coming from the trainer agent for these expert pairs.
Regarding the normalizations, we trained Behavioral Cloning (BC) and RILe across various data leakage levels, selecting the highest-scoring run (0% leakage RILe) as the baseline. Other scores and convergence steps are normalized by dividing by the score and convergence steps of the baseline (0% leakage RILe). For IQLearn, we used their reported numbers in their paper, as we couldn’t replicate their results with their code and hyperparameters.
Appendix E Additional Experiments
E.1 Robustness to Noise in the Expert Data